OpenStudy (anonymous):

3. Determine the zeros of: f(x) = x4 – x3 + 7x2 – 9x – 18.

OpenStudy (anonymous):

x=-1,2,-3i,3i

OpenStudy (anonymous):

i already did it, but im left with 2 and -1+- sqrt-35/-2??? what is wrong?

OpenStudy (anonymous):

It can be factored as (x-2)(x+1)(x^2+9)

OpenStudy (primeralph):

Easy way to know that 1 or -1 is a s zero is by seeing if all the coefficients somehow add up to 0. In this case, 1+1 +7+9 - 18 = 0. So -1 or 1 is a root.

OpenStudy (anonymous):

ok but look im supposed to do like this: Step 1: Use the Rational Root Theorem Step 2: Use Descartes' Rule of Signs Step 3: Use the Fundamental Theorem of Algebra Step 4: Test the factors Step 5: Write the depressed equation Step 6: Find the zeros Step 7: Look at the graph the function Step 8: Confirm the zeros by substituting

OpenStudy (anonymous):

and im left in step 6 with 2 and -1+-sqrt-35/-2 by using the quadratic formula! because of the discriminant!

OpenStudy (anonymous):

here's the table!

OpenStudy (anonymous):

jeezus christ you shouldn't be memorizing steps

OpenStudy (anonymous):

Step 1, wont even work if the constant term isn't rational.

OpenStudy (anonymous):

Also how does the fundamental theorem of algebra help you solve anything? Its an existence theorem not a constructive theorem.

OpenStudy (anonymous):

In addition the proofs of these techniques are pretty advanced, I wouldn't be inclined to use techniques you can't understand, but thats just me.

OpenStudy (anonymous):

i know right! but they want us to do things this way! I dont know what to do!!??

OpenStudy (anonymous):

i could factorize, i know but that;s out of " the steps"

OpenStudy (anonymous):

and they take points out :(

OpenStudy (anonymous):

So what do you need help with factoring, or finding the roots

OpenStudy (anonymous):

Do you actually have to learn the mathematics, or do you just want to get it done?

OpenStudy (anonymous):

Jack17 (x-2)(x+1)(x^2+9) is great and I know x^2+9 will result in x+-3i no, Iw ant to learn the mathematics, no one spends his whiole summer learning!!

OpenStudy (anonymous):

i want to understand thus thing if not i would just have copied from yahoo or here!

OpenStudy (anonymous):

sooo????

OpenStudy (mww):

I would try to find rational roots if any, then use polynomial division progressively with Factor theorem. This also helps to narrow down the possible roots.

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):



OpenStudy (anonymous):

now, is there any way it can be solved by completing the square or using the quadratic formula?

OpenStudy (anonymous):

because that's the only thing i can turn in in my assignment, base on the discriminant i got!

OpenStudy (anonymous):

which is -35 is that right?

OpenStudy (mww):

Another helpful tip is roots come in complex conjugate pairs if the coefficients of the polynomial are real. Completing the square or quad. formula only works once you narrow by poly division into quadratic factors. Sometimes you get irreducible quadratic factors and that means there are non-real complex roots.

OpenStudy (anonymous):

so if i get \[-1\pm \sqrt{-35}/ -2\]

OpenStudy (anonymous):

that means it has non real complex roots?

OpenStudy (mww):

is this for the same question? if you have a negative number in the discriminant then it has non real complex roots. You would rewrite it as sqrt(35) * i

OpenStudy (anonymous):

thta's what i was looking for, so it is\[-1+\sqrt{35*i}/-2\]

OpenStudy (mww):

um the i never goes in the sqrt! keep the i outside of it. You could write it as i sqrt(35) if you like. But that's not the answer for the question you posted. A different question?

OpenStudy (anonymous):

Determine the zeros of: f(x) = x4 – x3 + 7x2 – 9x – 18 <----- right?

OpenStudy (anonymous):

can u do it by completing the square please?

OpenStudy (anonymous):

sorry the quadratic formula, because i did it and that's what i got!

OpenStudy (mww):

Yep I might take you through this step by step: First test any rational roots. 1 and -1 are easy P(1) = (1)^4 - (1)^3 +7(1)^2 -9(1) - 18 which is not 0. P(-1) =... = 0 so x + 1 must be a factor by factor Theorem.

OpenStudy (mww):

Dividing the original poly by x+1 yields x^3 -2x^2 -9x - 18. Now we need to look for more roots. 1 did not work before, so try other factors of -18, say 2. P(2) also yields 0 so x + 2 must also be a root.

OpenStudy (mww):

Sorry x - 2 is a root not x + 2

OpenStudy (mww):

x - 2 being a factor, dividing x^3 -2x^2 -9x - 18 by this now gives x^2 + 9

OpenStudy (mww):

So solutions to x^2 + 9 = 0 gives you your complex roots. (x^2 + 9) = (x - 3i)(x + 3i) Our roots are then -1,2, 3i and -3i. There was no need to use quad. formula or complete the square. If you did use quadratic formula with x^2 + 9 = 0, you'd get b^2 - 4ac being (0)^2 - 4(1)(9) = -36 or (6i)^2. Then dividing by a = 2, gives you plus or minus of 3i

OpenStudy (anonymous):

\[-(b)\pm \sqrt{b ^{2}-4ac}/2(a)\] <------------ quadratic formula

OpenStudy (mww):

whoops yes i meant a = 1, 2a = 2

OpenStudy (anonymous):

whattt?